Hello, I am just wondering if someone could please please help find the solution to my problem;
the question is: A manager from a candy store mixed two different kinds of candy together, mint candy and cinnamon candy. The cinnamon candy cost 2,40/kg.
The mint candy cost 1,80/kg. The new mix cost 2,16$/kg. The manager makes 50kg
of the new mix. Determine the quantity of each type of candy that is necessary to be used in the new mixed.
Thank you so much for your help. i really appreciate it.
-kimberley

Answer by stanbon(75887)

The cinnamon candy cost 2,40/kg.
The mint candy cost 1,80/kg. The new mix cost 2,16$/kg. The manager makes 50kg
of the new mix. Determine the quantity of each type of candy that is necessary to be used in the new mixed.
---------------
Let amount of cin candy be "x" ; Value of this is 2.40x dollars
Amount of mint candy is "50-x" ; Value of this is 1.80(50-x)=90-1.8x dollars
Amount of mix candy is 50 kg ; Value of this is 2.16(50)=108 dollars
EQUATION:
value + value = value
2.40x + 90-1.8x=108
0.6x=18
x=30 kg (amount of cinnamon candy in mix)
50-x=20 kn (amount of mint candy in mix)
Cheers,
Stan H.